A tumor is injected with 0.8 grams of Iodine-125, which has a decay rate of 1.15% per day. Write an exponential model representing the amount of Iodine-125 remaining in the tumor after t days. Then use the formula to find the amount of Iodine-125 that would remain in the tumor after 60 days.
1
Expert's answer
2020-09-21T15:53:48-0400

 we can use the model for continuous exponential decay

"m = m_0e^{kt}"

where m0 =0.8 initial mass

t = 60 time in days

k = -1.15/100 = -0.0115 continuous growth rate

e base of natural logarithm

m ending value mass


"m = m_0e^{kt} = 0.8*e^{-0.0115*60} = 0.8*e^{-0.69} \approx 0.401"

Answer: 0.401g